a 12 resistor and a 15 resistor are connected in series Across 9.0 V potential difference what is the current in the circuit


1
Expert's answer
2021-05-24T15:48:55-0400

Gives

"R_1=12
\Omega"

R2=15"\Omega"

v =9.0V

Series combination

"R=R_1+R_2"

"R=12+15\Omega"

We know that

"V=IR"

"I=\frac{V}{R}"

Put value

"I=\frac{9.0}{27}=0.33A"